OpenStudy (anonymous):

Solve using the addition prinsiple. Don't forget to check. - 2/3 + y = - 3/4

hero (hero):

Hint: Add 2/3 to both sides

OpenStudy (anonymous):

You mean - 2/3 + y = -3/4 + 2/3?

OpenStudy (anonymous):

to check plug in the resulting value for y and see if it turns out true :)

hero (hero):

If you add 2/3 to both sides you end up with -2/3 + 2/3 + y = -3/4 + 2/3

hero (hero):

-2/3 + 2/3 = 0 leaving you with y = -3/4 + 2/3

hero (hero):

Now all you have to do is simplify the right side by adding fractions

OpenStudy (anonymous):

ok thanks